# Optimal Methods for Unknown Piecewise Smooth Problems I: Convex Optimization

### v1

@@ 8.2 Two-stage stochastic linear programming
Now, we have proven that there exists an algorithm $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}_{\text{zr}}$ and matrix $U\in\mathbb{R}^{d\times d}$ satisfying the first claim of Lemma A.1. In the following, we show that the algorithm $\mathcal{Z}_{\mathcal{A}}$ and matrix $U$ has property (64) for any $\varepsilon$. If $T_{\varepsilon}\left(\mathcal{A},f_{U}\right)>d$, we are done. Otherwise $T_{\varepsilon}\left(\mathcal{A},f_{U}\right)\leq d$ and we have
$\begin{split}&T_{\varepsilon}\left(\mathcal{A},f_{U}\right):=T_{\varepsilon}\left(\mathcal{A}\left[f_{U}\right],f_{U}\right)\\ &=\inf\left\{t\in\mathbb{N}:f_{U}\left(a^{t}\right)-f_{U}\left(a^{*}\right)\leq\varepsilon\right\}\\ &\overset{(a)}{=}\inf\left\{t\in\mathbb{N}:f\left(U^{\top}a^{t}\right)-f^{*}\leq\varepsilon\right\}=T_{\varepsilon}\left(U^{\top}\mathcal{A}\left[f_{U}\right],f\right)\\ &=T_{\varepsilon}\left(U^{\top}\mathcal{A}\left[f_{U}\right],f\right)\overset{(b)}{=}T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\end{split}$ (69)
@@
$T_{\varepsilon}\left(\mathcal{A},f_{U\left(f,\mathcal{A}\right)}\right)\geq\min\left\{d,T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\right\}$ (70)
@@
$\mathcal{F}_{d}^{\tilde{d}}=\left\{f_{U\left(f,\mathcal{A}\right)}:f_{U\left(f,\mathcal{A}\right)}\in\mathcal{F}^{\tilde{d}},f_{U\left(f,\mathcal{A}\right)}=f\left(U\left(f,\mathcal{A}\right)^{\top}x\right),U\left(f,\mathcal{A}\right)\in\mathbb{R}^{\tilde{d}\times d},U\left(f,\mathcal{A}\right)^{\top}U\left(f,\mathcal{A}\right)=I_{d}\right\}$ (71)
@@
$\begin{split}&\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{det}},\mathcal{F}^{\tilde{d}}\right)\\ &=\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{\tilde{f}\in\mathcal{F}^{\tilde{d}}}T_{\varepsilon}\left(\mathcal{A},\tilde{f}\right)\overset{(a)}{\geq}\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{f_{U}\in\mathcal{F}_{d}^{\tilde{d}}\subseteq\mathcal{F}^{\tilde{d}}}T_{\varepsilon}\left(\mathcal{A},f_{U}\right)\overset{(b)}{\geq}\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{f_{U}\in\mathcal{F}_{d}^{\tilde{d}}\subseteq\mathcal{F}^{\tilde{d}}}\min\left\{d,T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\right\}\\ &\overset{(c)}{=}\min\left\{d,\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{f\in\mathcal{F}^{d}}T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\right\}\overset{(d)}{\geq}\min\left\{d,\inf_{\mathcal{Z}\in\mathbb{A}_{\text{zr}}}\sup_{f\in\mathcal{F}^{d}}T_{\varepsilon}\left(\mathcal{Z},f\right)\right\}=\min\left\{d,\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{zr}},\mathcal{F}^{d}\right)\right\}\end{split}\ ,$ (72)
where $(a)$ holds since the feasible region $\mathcal{F}_{d}^{\tilde{d}}\subseteq\mathcal{F}^{\tilde{d}}$ is a subset of $\mathcal{F}^{\tilde{d}}$ (parameterized by matrix $U$ and function $f$), $(b)$ follows from (70), $(c)$ holds by replacing the optimizer from $f_{U}$ to $f$ and $(d)$ is due to $\mathcal{A}\in\mathbb{A}_{\det}$ then the feasible region of $\mathcal{Z}_{\mathcal{A}}$ is a subset of $\mathbb{A}_{\text{zr}}$. As we chose $d$ satisfying $\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{zr}},\mathcal{F}^{d}\right)<d$, hence, inequality (72) implies that $\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{det}},\mathcal{F}^{\tilde{d}}\right)\geq\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{zr}},\mathcal{F}^{d}\right)$. ∎∎
@@ Appendix B How to optimize the subproblem
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}\left\|x-\bar{y}\right\|^{2}\ ,\ \operatornamewithlimits{s.t.}Ax\leq b\ ,$ (73)
@@
An important observation is that the dual of (73) is
$\min_{\lambda\geq 0}\ \ \frac{1}{2}\lambda^{\top}(AA^{\top})\lambda+\lambda^{\top}\left(b-A\bar{y}\right)\ ,$ (74)
@@
$\min_{\lambda\geq 0}\ \ \frac{1}{2}\left\|A^{\top}\lambda-(\bar{y}-\tilde{b})\right\|^{2}\ ,$ (75)

### v2

@@ 8.2 Two-stage stochastic linear programming
Now, we have proven that there exists an algorithm $\mathcal{Z}_{\mathcal{A}}\in\mathbb{A}_{\text{zr}}$ and matrix $U\in\mathbb{R}^{d\times d}$ satisfying the first claim of Lemma A.1. In the following, we show that the algorithm $\mathcal{Z}_{\mathcal{A}}$ and matrix $U$ has property (A.1) for any $\varepsilon$. If $T_{\varepsilon}\left(\mathcal{A},f_{U}\right)>d$, we are done. Otherwise $T_{\varepsilon}\left(\mathcal{A},f_{U}\right)\leq d$ and we have
$\begin{split}&T_{\varepsilon}\left(\mathcal{A},f_{U}\right):=T_{\varepsilon}\left(\mathcal{A}\left[f_{U}\right],f_{U}\right)\\ &=\inf\left\{t\in\mathbb{N}:f_{U}\left(a^{t}\right)-f_{U}\left(a^{*}\right)\leq\varepsilon\right\}\\ &\overset{(a)}{=}\inf\left\{t\in\mathbb{N}:f\left(U^{\top}a^{t}\right)-f^{*}\leq\varepsilon\right\}=T_{\varepsilon}\left(U^{\top}\mathcal{A}\left[f_{U}\right],f\right)\\ &=T_{\varepsilon}\left(U^{\top}\mathcal{A}\left[f_{U}\right],f\right)\overset{(b)}{=}T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\end{split}$ (A.6)
@@
$T_{\varepsilon}\left(\mathcal{A},f_{U\left(f,\mathcal{A}\right)}\right)\geq\min\left\{d,T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\right\}$ (A.7)
@@
$\mathcal{F}_{d}^{\tilde{d}}=\left\{f_{U\left(f,\mathcal{A}\right)}:f_{U\left(f,\mathcal{A}\right)}\in\mathcal{F}^{\tilde{d}},f_{U\left(f,\mathcal{A}\right)}=f\left(U\left(f,\mathcal{A}\right)^{\top}x\right),U\left(f,\mathcal{A}\right)\in\mathbb{R}^{\tilde{d}\times d},U\left(f,\mathcal{A}\right)^{\top}U\left(f,\mathcal{A}\right)=I_{d}\right\}$ (A.8)
@@
$\begin{split}&\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{det}},\mathcal{F}^{\tilde{d}}\right)\\ &=\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{\tilde{f}\in\mathcal{F}^{\tilde{d}}}T_{\varepsilon}\left(\mathcal{A},\tilde{f}\right)\overset{(a)}{\geq}\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{f_{U}\in\mathcal{F}_{d}^{\tilde{d}}\subseteq\mathcal{F}^{\tilde{d}}}T_{\varepsilon}\left(\mathcal{A},f_{U}\right)\overset{(b)}{\geq}\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{f_{U}\in\mathcal{F}_{d}^{\tilde{d}}\subseteq\mathcal{F}^{\tilde{d}}}\min\left\{d,T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\right\}\\ &\overset{(c)}{=}\min\left\{d,\inf_{\mathcal{A}\in\mathbb{A}_{\text{det}}}\sup_{f\in\mathcal{F}^{d}}T_{\varepsilon}\left(\mathcal{Z}_{\mathcal{A}},f\right)\right\}\overset{(d)}{\geq}\min\left\{d,\inf_{\mathcal{Z}\in\mathbb{A}_{\text{zr}}}\sup_{f\in\mathcal{F}^{d}}T_{\varepsilon}\left(\mathcal{Z},f\right)\right\}=\min\left\{d,\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{zr}},\mathcal{F}^{d}\right)\right\}\end{split}\ ,$ (A.9)
where $(a)$ holds since the feasible region $\mathcal{F}_{d}^{\tilde{d}}\subseteq\mathcal{F}^{\tilde{d}}$ is a subset of $\mathcal{F}^{\tilde{d}}$ (parameterized by matrix $U$ and function $f$), $(b)$ follows from (A.7), $(c)$ holds by replacing the optimizer from $f_{U}$ to $f$ and $(d)$ is due to $\mathcal{A}\in\mathbb{A}_{\det}$ then the feasible region of $\mathcal{Z}_{\mathcal{A}}$ is a subset of $\mathbb{A}_{\text{zr}}$. As we chose $d$ satisfying $\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{zr}},\mathcal{F}^{d}\right)<d$, hence, inequality (A.9) implies that $\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{det}},\mathcal{F}^{\tilde{d}}\right)\geq\mathcal{T}_{\varepsilon}\left(\mathbb{A}_{\text{zr}},\mathcal{F}^{d}\right)$. ∎∎
@@ Appendix B How to optimize the subproblem
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}\left\|x-\bar{y}\right\|^{2}\ ,\ \operatornamewithlimits{s.t.}Ax\leq b\ ,$ (B.1)
@@
An important observation is that the dual of (B.1) is
$\min_{\lambda\geq 0}\ \ \frac{1}{2}\lambda^{\top}(AA^{\top})\lambda+\lambda^{\top}\left(b-A\bar{y}\right)\ ,$ (B.2)
@@
$\min_{\lambda\geq 0}\ \ \frac{1}{2}\left\|A^{\top}\lambda-(\bar{y}-\tilde{b})\right\|^{2}\ ,$ (B.3)